# Crosscap states and duality of Ising field theory in two dimensions

### v1

@@ Abstract
which converges quickly as a power series in $\eta$. The expansion $G_{\pm}^{\sigma}(\eta)=(1-\eta)^{-2h_{\sigma}}(\sum_{n=0}^{\infty}a_{n}\eta^{n}\pm\eta^{h_{\varepsilon}}\sum_{n=0}^{\infty}b_{n}\eta^{n})$ is precisely the conformal partial wave decomposition [32, 33], consistent with the fusion rule $\sigma\sigma\sim 1+\varepsilon$. By choosing a suitable cutoff ($n_{\mathrm{max}}\approx 40$) to truncate the contributions from the descendant fields with large conformal weights, we can perform the numerical integration in Eq. (S178) with very high precision. The final results are given by
@@
F.4 D. A simple derivation of the one- and two-point crosscap correlators
For completeness, we provide a self-contained derivation of the one- and two-point crosscap correlators [Eqs. (S162) and (S163)] based on the “doubling trick” in the boundary CFT [65].
We start with the crosscap state expressed as the combination of the crosscap Ishibashi states: $|\mathcal{C}\rangle=\sum_{a}\mathcal{A}_{a}|a\rangle\rangle_{\mathcal{C}}$. The one-point crosscap correlator of the primary field $\varphi$ with conformal weight $h_{a}=\bar{h}_{a}$ is given by:
$\langle\psi_{0}(0)|\varphi(w,\bar{w})|\mathcal{C}\rangle=\mathcal{A}_{a}\langle\psi_{0}(0)|\varphi(w,\bar{w})|a\rangle\rangle_{\mathcal{C}}\,.$ (S187)
The key observation is that the crosscap Ishibashi state $|a\rangle\rangle_{\mathcal{C}}$ can be related to the ordinary Ishibashi state $|a\rangle\rangle$ as [see Eq. (S67)]:
$|a\rangle\rangle_{\mathcal{C}}=(-1)^{L_{0}-h_{a}}|a\rangle\rangle\,,$ (S188)
from which we obtain
$\langle\psi_{0}(0)|\varphi_{a}(w,\bar{w})|a\rangle\rangle_{\mathcal{C}}$ $=\langle\psi_{0}(0)|\varphi_{a}(w,\bar{w})e^{i\pi(L_{0}-h_{a})}|a\rangle\rangle$
$=e^{-i\pi h_{a}}\langle\psi_{0}(0)|e^{-i\pi L_{0}}\varphi_{a}(w,\bar{w})e^{i\pi L_{0}}|a\rangle\rangle$
$=e^{-i\pi h_{a}}\langle\psi_{0}(0)|\varphi_{a}(-w,\bar{w})|a\rangle\rangle\,,$ (S189)
where we used ${w^{\prime}}^{L_{0}}\varphi(w,\bar{w}){w^{\prime}}^{-L_{0}}=\varphi(ww^{\prime},\bar{w})$, with $w^{\prime}=e^{-i\pi}$, in the radial quantization picture.
We further consider the Möbius transformation
$\zeta=S(w)=\frac{i(w-i)}{w+i}\,,$ (S190)
which maps the $\mathbb{RP}^{2}$ to the upper half plane $\mathbb{HP}_{+}$. Consequently, the primary field $\varphi_{a}$ transforms as
$\varphi_{a}(w,\bar{w})=\left[S^{\prime}(w)\bar{S}^{\prime}(\bar{w})\right]^{h_{a}}\varphi_{a}(\zeta,\bar{\zeta})=\left[\frac{2}{(w+i)^{2}}\cdot\frac{2}{(\bar{w}-i)^{2}}\right]^{h_{a}}\varphi_{a}(\zeta,\bar{\zeta})\,.$ (S191)
We can imagine that the half plane has an associated boundary state, denoted as $|\mathcal{B}\rangle=\sum_{a}\tilde{\mathcal{A}}_{a}|a\rangle\rangle$, then the one-point boundary correlator reads
$\langle\psi_{0}(0)|\varphi_{a}(\zeta,\bar{\zeta})|\mathcal{B}\rangle=\tilde{\mathcal{A}}_{a}\langle\psi_{0}(0)|\varphi_{a}(\zeta,\bar{\zeta})|a\rangle\rangle\,.~{}$ (S192)
On the other hand, using the “doubling trick”, the one-point boundary correlator on the half plane is equal to the two-point chiral correlator on the $\zeta$-plane:
$\langle\psi_{0}(0)|\varphi_{a}(\zeta,\bar{\zeta})|\mathcal{B}\rangle=\langle\phi_{a}(\zeta)\phi_{a}(\bar{\zeta})\rangle=\frac{\tilde{\mathcal{A}}_{a}}{(\zeta-\bar{\zeta})^{2h_{a}}}\,,$ (S193)
where $\phi_{a}$ is the chiral part of the primary field: $\varphi_{a}(\zeta,\bar{\zeta})$. The amplitude $\tilde{\mathcal{A}}_{a}$ is fixed via the OPE: $\phi_{a}\phi_{b}\sim\delta_{a,b}+\cdots$. Comparing Eq. (S192) and Eq. (S193), we determine the Ishibashi boundary correlator as
$\langle\psi_{0}(0)|\varphi_{a}(\zeta,\bar{\zeta})|a\rangle\rangle=\frac{1}{(\zeta-\bar{\zeta})^{2h_{a}}}\,.$ (S194)
Therefore, the one-point crosscap correlator is obtained via the conformal transformation $\zeta=S(w)$,
$\langle\psi_{0}(0)|\varphi_{a}(w,\bar{w})|\mathcal{C}\rangle$ $=\mathcal{A}_{a}e^{-i\pi h_{a}}\langle\psi_{0}(0)|\varphi_{a}(-w,\bar{w})|a\rangle\rangle$
@@
$=\mathcal{A}_{a}e^{-i\pi h_{a}}\left[S^{\prime}(-w)\bar{S}^{\prime}(\bar{w})\right]^{h_{a}}\langle\psi_{0}(0)|\varphi_{a}(-\frac{1}{\zeta},\bar{\zeta})|a\rangle\rangle$
$=\mathcal{A}_{a}e^{-i\pi h_{a}}\left[\frac{2}{(w-i)(\bar{w}-i)}\right]^{2h_{a}}\frac{1}{(-\frac{1}{\zeta}-\bar{\zeta})^{2h_{a}}}$
$=\frac{\mathcal{A}_{a}}{(1+|w|^{2})^{2h_{a}}}\,,$ (S195)
where $\zeta=S(w)$ and $-1/\zeta=S(-w)$.
The two-point crosscap correlator should behave as
$\langle 0|\varphi_{a}(w_{1},\bar{w}_{1})\varphi_{a}(w_{2},\bar{w}_{2})|\mathcal{C}\rangle=\frac{\mathcal{A}_{1}}{|w_{1}-w_{2}|^{4h_{a}}}\cdot G_{a}(w_{1},\bar{w}_{1},w_{2},\bar{w}_{2})\,,$ (S196)
via the OPE: $\varphi_{a}\varphi_{a}\sim 1+\cdots$. The dimensionless prefactor $G_{a}$ remains to be fixed.
Using the similar argument as the one-point function case, the two-point crosscap correlator should be related to the four-point chiral correlator on the $\zeta$-plane
$\langle\phi_{a}(-\frac{1}{\zeta_{1}})\phi_{a}(\bar{\zeta}_{1})\phi_{a}(-\frac{1}{\zeta_{2}})\phi_{a}(\bar{\zeta}_{2})\rangle$ (S197)
via the “doubling trick”. Therefore, the undetermined prefactor $G_{a}$ must be a function of the cross ratio $\eta$ on the $\zeta$-plane:
$\eta=\frac{(-\frac{1}{\zeta_{1}}+\frac{1}{\zeta_{2}})(\bar{\zeta}_{1}-\bar{\zeta}_{2})}{(-\frac{1}{\zeta_{1}}-\bar{\zeta}_{1})(-\frac{1}{\zeta_{2}}-\bar{\zeta}_{2})}=\frac{|\zeta_{1}-\zeta_{2}|^{2}}{(1+|\zeta_{1}|^{2})(1+|\zeta_{2}|^{2})}=\frac{|w_{1}-w_{2}|^{2}}{(1+|w_{1}|^{2})(1+|w_{2}|^{2})}\,,$ (S198)
where in the last equality, we used the fact that the cross ratio $\eta$ is invariant under the Möbius transformation: $\zeta=S(w)$.
In the above derivation, we only used the Möbius transformation, which is the global conformal transformation in 2D spacetime. Therefore, this approach can be generalized to higher dimensions [32, 33].

### v2

@@ Abstract
which converges quickly as a power series in $\eta$. The expansion $G_{\pm}^{\sigma}(\eta)=(1-\eta)^{-2h_{\sigma}}(\sum_{n=0}^{\infty}a_{n}\eta^{n}\pm\eta^{h_{\varepsilon}}\sum_{n=0}^{\infty}b_{n}\eta^{n})$ is precisely the conformal partial wave decomposition [33, 34], consistent with the fusion rule $\sigma\times\sigma\sim\mathbbm{1}+\varepsilon$. By choosing a suitable cutoff ($n_{\mathrm{max}}\approx 40$) to truncate the contributions from the descendant fields with large conformal weights, we can perform the numerical integration in Eq. (S197) with very high precision. The final results are given by
@@
F.5 E. $\mathbb{Z}_{3}$ parafermion CFT with thermal perturbation
For the $\mathbb{Z}_{3}$ parafermion CFT with thermal perturbation (perturbation operator has conformal weight $h_{\varepsilon}=\bar{h}_{\varepsilon}=2/5$), the relevant crosscap coefficients of the standard PSS crosscap state are $\mathcal{A}_{\mathbbm{1}}=(3+6/\sqrt{5})^{1/4}$ and $\mathcal{A}_{\varepsilon}=(3-6/\sqrt{5})^{1/4}$, which gives the following first-order perturbation correction to the crosscap overlap:
@@
$\langle\psi_{0}(s)|\mathcal{C}_{\pm}^{\mathbb{Z}_{3}}\rangle=\left(3+\frac{6}{\sqrt{5}}\right)^{1/4}\pm\frac{5}{8(2\pi)^{1/5}}\left(3-\frac{6}{\sqrt{5}}\right)^{1/4}\,\frac{\Gamma(2/5)\Gamma(7/5)}{\Gamma(4/5)}\cdot s+\mathcal{O}(s^{2})=1.54401\pm 0.54881\cdot s+\mathcal{O}(s^{2})\,$, (S210)
Leading-order expansion coefficients of the crosscap overlap are expected to be useful in identifying critical theories in lattice models simulations. As an example, we consider the three-state quantum clock chain
with
$\sigma=\begin{pmatrix}0&1&0\\ 0&0&1\\ 1&0&0\end{pmatrix},\quad\tau=\begin{pmatrix}1&0&0\\ 0&e^{2\pi\mathrm{i}/3}&0\\ 0&0&e^{4\pi\mathrm{i}/3}\end{pmatrix}$ (S212)
being $\mathbb{Z}_{3}$ spin matrices, where periodic boundary condition is imposed. At the critical point $h_{3}=1$, the model (S211) is described by the $\mathbb{Z}_{3}$ parafermion CFT, with velocity $v=\frac{3\sqrt{3}}{2}$ determined from the exact solution [69]. Away from (but near) $h_{3}=1$, the model is ordered (with spontaneous $\mathbb{Z}_{3}$ symmetry broken) for $h_{3}<1$ and disordered for $h_{3}>1$, whose low-energy effective theory is just the $\mathbb{Z}_{3}$ parafermion CFT with thermal perturbation. The dimensionless coupling $s$ for the lattice model (S211) is given by $s=\frac{\sqrt{\mathcal{N}_{\varepsilon}}}{v}(1-h_{3})(vN)^{6/5}$ [43], where $\mathcal{N}_{\varepsilon}=\lim_{r\rightarrow\infty}\lim_{N\rightarrow\infty}r^{8/5}\langle(\tau^{{\dagger}}_{j}+\tau_{j})(\tau^{{\dagger}}_{j+r}+\tau_{j+r})\rangle_{\mathrm{c}}=0.315$ ($\langle\cdots\rangle_{\mathrm{c}}$: connected correlator evaluated at the critical point $h_{3}=1$) is the normalization of the perturbation operator on the lattice. The crosscap overlap between the ground state of Eq. (S211) and the lattice crosscap state $|\mathcal{C}_{\mathrm{latt}}\rangle=\prod_{j=1}^{N/2}\left(\sum_{\alpha=1}^{3}|\alpha\rangle_{j}|\alpha\rangle_{j+N/2}\right)$ has been calculated using DMRG. The numerical results shown in Fig. S2 are in good agreement with the conformal perturbation theory result in Eq. (S210).
Figure S2: Crosscap overlap of the three-state clock model compared with the leading-order (1st order) conformal perturbation prediction. Numerical results with three different chain lengths ($N=20,30,40$) are shown with different colored symbols. For all three data sets, fitting their slopes gives (almost) the same numerical value $\mathcal{C}^{\mathrm{num}}_{1}\approx 0.55$.